OpenStudy (anonymous):

\[2^{x} \times 4^{x} \times 8^{x+1} = 2^{21}\]

OpenStudy (whpalmer4):

Remember some properties of exponents: \[(ab)^n = a^nb^n\]\[a^n*a^m = a^{n+m}\]

OpenStudy (mathstudent55):

The entire left side can be written as a product of powers with 2 as base. Since they are all multiplied together, keep base of 2 and add the exponents. Then if a^m = a^n, it must be that m = n.